Grade 11Mechanics

two straight wire the length of wire id 10cm the distance between 0.5cm of palleral wire,the sheet makes of water, if the distance between wire increased 1mm of wire, so find the work done? surface tension of water is 0.072N/m

sarthak bhardwaj
7 Years agoGrade 11

1 Answer

Kapil Khare
7 Years ago
We have,
Length of wire = 10cm = 0.1m
Change in distance between the two wires = 1mm = 0.001m
So, change in area of one water sheet between the wires = 0.1*0.001 = 10-4m2
Change in Surface energy = (surface tension)*(change in area)
Now there are two surfaces of water. If we consider the wires to be in the plane of page, then one surface will be towards the reader and another will be on opposite side that is away from reader.
So, total change in area of water sheet = 2*10-4m2 
\impliesChange in Surface energy = (0.072)*(2*10-4) = 1.44*10-3Joules
The total work done to increase the distance between the wires will store as surface energy that means total work done is change in suface energy.
\implies Total work done in increasing the distance between the wires = 1.44*10-3Joules